\begin{thebibliography}{100}

\bibitem{GW170817-GRB170817A}
B.~Abbott, R.~Abbott, T.~Abbott, F.~Acernese, K.~Ackley, C.~Adams, T.~Adams,
  P.~Addesso, R.~Adhikari, V.~Adya, et~al.
\newblock Gravitational waves and gamma-rays from a binary neutron star merger:
  Gw170817 and grb 170817a.
\newblock {\em The Astrophysical Journal Letters}, 848(2):L13, 2017.

\bibitem{GW170817_LVC_progenitor}
B.~Abbott, R.~Abbott, T.~Abbott, F.~Acernese, K.~Ackley, C.~Adams, T.~Adams,
  P.~Addesso, R.~Adhikari, V.~Adya, et~al.
\newblock On the progenitor of binary neutron star merger gw170817.
\newblock {\em The Astrophysical Journal Letters}, 850(2):L40, 2017.

\bibitem{LVCK_prospects}
B.~P. Abbott, R.~Abbott, T.~Abbott, M.~Abernathy, F.~Acernese, K.~Ackley,
  C.~Adams, T.~Adams, P.~Addesso, R.~Adhikari, et~al.
\newblock Prospects for observing and localizing gravitational-wave transients
  with advanced ligo, advanced virgo and kagra.
\newblock {\em Living Reviews in Relativity}, 21(1):3, 2018.

\bibitem{GW_IFO_3rd_Gen_US}
B.~P. Abbott, R.~Abbott, T.~Abbott, M.~Abernathy, K.~Ackley, C.~Adams,
  P.~Addesso, R.~Adhikari, V.~Adya, C.~Affeldt, et~al.
\newblock Exploring the sensitivity of next generation gravitational wave
  detectors.
\newblock {\em Classical and Quantum Gravity}, 34(4):044001, 2017.

\bibitem{GW170817_LVC}
B.~P. Abbott, R.~Abbott, T.~Abbott, F.~Acernese, K.~Ackley, C.~Adams, T.~Adams,
  P.~Addesso, R.~Adhikari, V.~Adya, et~al.
\newblock Gw170817: observation of gravitational waves from a binary neutron
  star inspiral.
\newblock {\em Physical Review Letters}, 119(16):161101, 2017.

\bibitem{LIV_090510_limit_2009}
A.~Abdo, M.~Ackermann, M.~Ajello, K.~Asano, W.~Atwood, M.~Axelsson, L.~Baldini,
  J.~Ballet, G.~Barbiellini, M.~Baring, et~al.
\newblock A limit on the variation of the speed of light arising from quantum
  gravity effects.
\newblock {\em Nature}, 462(7271):331, 2009.

\bibitem{GW170817_afterglow_latettime_2018}
K.~Alexander, R.~Margutti, P.~Blanchard, W.~Fong, E.~Berger, A.~Hajela,
  T.~Eftekhari, R.~Chornock, P.~Cowperthwaite, D.~Giannios, et~al.
\newblock A decline in the x-ray through radio emission from gw170817 continues
  to support an off-axis structured jet.
\newblock {\em arXiv preprint arXiv:1805.02870}, 2018.

\bibitem{GW170817_kilonova_models_early}
I.~Arcavi.
\newblock The first hours of the gw170817 kilonova and the importance of early
  optical and ultraviolet observations for constraining emission models.
\newblock {\em The Astrophysical Journal Letters}, 855(2):L23, 2018.

\bibitem{CTA_2}
I.~{Bartos}, T.~{Di Girolamo}, J.~R. {Gair}, M.~{Hendry}, I.~S. {Heng}, T.~B.
  {Humensky}, S.~{M{\'a}rka}, Z.~{M{\'a}rka}, C.~{Messenger}, R.~{Mukherjee},
  D.~{Nieto}, P.~{O'Brien}, and M.~{Santander}.
\newblock {Strategies for the follow-up of gravitational wave transients with
  the Cherenkov Telescope Array}.
\newblock {\em \mnras}, 477:639--647, June 2018.

\bibitem{CTA_1}
I.~{Bartos}, P.~{Veres}, D.~{Nieto}, V.~{Connaughton}, B.~{Humensky},
  K.~{Hurley}, S.~{M{\'a}rka}, P.~{M{\'e}sz{\'a}ros}, R.~{Mukherjee},
  P.~{O'Brien}, and J.~P. {Osborne}.
\newblock {Cherenkov Telescope Array is well suited to follow up
  gravitational-wave transients}.
\newblock {\em \mnras}, 443:738--749, Sept. 2014.

\bibitem{NS_EOS_bauswein_2018}
A.~Bauswein, N.-U.~F. Bastian, D.~B. Blaschke, K.~Chatziioannou, J.~A. Clark,
  T.~Fischer, and M.~Oertel.
\newblock Identifying a first-order phase transition in neutron star mergers
  through gravitational waves.
\newblock {\em arXiv preprint arXiv:1809.01116}, 2018.

\bibitem{GRBs_photospheric}
A.~Beloborodov and P.~M{\'e}sz{\'a}ros.
\newblock Photospheric emission of gamma-ray bursts.
\newblock {\em Space Science Reviews}, 207(1-4):87--110, 2017.

\bibitem{2015CQGra..32x3001B}
E.~Berti, E.~Barausse, V.~Cardoso, L.~Gualtieri, P.~Pani, U.~Sperhake, L.~C.
  Stein, N.~Wex, K.~Yagi, T.~Baker, et~al.
\newblock Testing general relativity with present and future astrophysical
  observations.
\newblock {\em Classical and Quantum Gravity}, 32(24):243001, 2015.

\bibitem{joint_blackburn}
L.~Blackburn, M.~S. Briggs, J.~Camp, N.~Christensen, V.~Connaughton, P.~Jenke,
  R.~A. Remillard, and J.~Veitch.
\newblock High-energy electromagnetic offline follow-up of ligo-virgo
  gravitational-wave binary coalescence candidate events.
\newblock {\em The Astrophysical Journal Supplement Series}, 217(1):8, 2015.

\bibitem{jets_blandford_znajek}
R.~D. Blandford and R.~L. Znajek.
\newblock Electromagnetic extraction of energy from kerr black holes.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  179(3):433--456, 1977.

\bibitem{Blinnikov1984}
S.~I. {Blinnikov}, I.~D. {Novikov}, T.~V. {Perevodchikova}, and A.~G.
  {Polnarev}.
\newblock {Exploding Neutron Stars in Close Binaries}.
\newblock {\em Soviet Astronomy Letters}, 10:177--179, Apr. 1984.

\bibitem{NSM_Bloom_1999}
J.~S. {Bloom}, S.~{Sigurdsson}, and O.~R. {Pols}.
\newblock {The spatial distribution of coalescing neutron star binaries:
  implications for gamma-ray bursts}.
\newblock {\em \mnras}, 305:763--769, May 1999.

\bibitem{Planck_2018_cosmo}
P.~Collaboration.
\newblock {Planck 2018 results. VI. Cosmological parameters}.
\newblock {\em ArXiv e-prints}, July 2018.

\bibitem{LIV_SME}
D.~Colladay and V.~A. Kosteleck{\`y}.
\newblock Lorentz-violating extension of the standard model.
\newblock {\em Physical Review D}, 58(11):116002, 1998.

\bibitem{rprocess_2_cote}
B.~C{\^o}t{\'e}, C.~L. Fryer, K.~Belczynski, O.~Korobkin,
  M.~Chru{\'s}li{\'n}ska, N.~Vassh, M.~R. Mumpower, J.~Lippuner, T.~M. Sprouse,
  R.~Surman, et~al.
\newblock The origin of r-process elements in the milky way.
\newblock {\em The Astrophysical Journal}, 855(2):99, 2018.

\bibitem{GW170817_kilonova_optical}
D.~Coulter, R.~Foley, C.~Kilpatrick, M.~Drout, A.~Piro, B.~Shappee, M.~Siebert,
  J.~Simon, N.~Ulloa, D.~Kasen, et~al.
\newblock Swope supernova survey 2017a (sss17a), the optical counterpart to a
  gravitational wave source.
\newblock {\em Science}, page eaap9811, 2017.

\bibitem{dalal_2006}
N.~Dalal, D.~E. Holz, S.~A. Hughes, and B.~Jain.
\newblock Short grb and binary black hole standard sirens as a probe of dark
  energy.
\newblock {\em Physical Review D}, 74(6):063006, 2006.

\bibitem{dominik_2013}
M.~Dominik, K.~Belczynski, C.~Fryer, D.~E. Holz, E.~Berti, T.~Bulik, I.~Mandel,
  and R.~O'Shaughnessy.
\newblock Double compact objects. ii. cosmological merger rates.
\newblock {\em The Astrophysical Journal}, 779(1):72, 2013.

\bibitem{rprocess_3_drout}
M.~Drout, A.~Piro, B.~Shappee, C.~Kilpatrick, J.~Simon, C.~Contreras,
  D.~Coulter, R.~Foley, M.~Siebert, N.~Morrell, et~al.
\newblock Light curves of the neutron star merger gw170817/sss17a: Implications
  for r-process nucleosynthesis.
\newblock {\em Science}, page eaaq0049, 2017.

\bibitem{GRB170817A_Eichler}
D.~Eichler.
\newblock Short gamma-ray bursts viewed from far off-axis.
\newblock {\em The Astrophysical Journal Letters}, 869(1):L4, 2018.

\bibitem{Eichler_1989}
D.~{Eichler}, M.~{Livio}, T.~{Piran}, and D.~N. {Schramm}.
\newblock {Nucleosynthesis, neutrino bursts and gamma-rays from coalescing
  neutron stars}.
\newblock {\em \nat}, 340:126--128, July 1989.

\bibitem{GW170817_kilonova_UV}
P.~Evans, S.~Cenko, J.~Kennea, S.~Emery, N.~Kuin, O.~Korobkin, R.~Wollaeger,
  C.~Fryer, K.~Madsen, F.~Harrison, et~al.
\newblock Swift and nustar observations of gw170817: Detection of a blue
  kilonova.
\newblock {\em Science}, 358(6370):1565--1570, 2017.

\bibitem{BNS_review_rasio}
J.~A. Faber and F.~A. Rasio.
\newblock Binary neutron star mergers.
\newblock {\em Living Reviews in Relativity}, 15(1):8, 2012.

\bibitem{SGRBs_fong}
W.-f. Fong, E.~Berger, R.~Margutti, and B.~A. Zauderer.
\newblock A decade of short-duration gamma-ray burst broadband afterglows:
  energetics, circumburst densities, and jet opening angles.
\newblock {\em The Astrophysical Journal}, 815(2):102, 2015.

\bibitem{NS_EOS_foucart_2012}
F.~Foucart.
\newblock Black-hole--neutron-star mergers: Disk mass predictions.
\newblock {\em Physical Review D}, 86(12):124007, 2012.

\bibitem{freedman2001final}
W.~L. Freedman, B.~F. Madore, B.~K. Gibson, L.~Ferrarese, D.~D. Kelson,
  S.~Sakai, J.~R. Mould, R.~C. Kennicutt~Jr, H.~C. Ford, J.~A. Graham, et~al.
\newblock Final results from the hubble space telescope key project to measure
  the hubble constant.
\newblock {\em The Astrophysical Journal}, 553(1):47, 2001.

\bibitem{Fryer_1999}
C.~L. {Fryer}, S.~E. {Woosley}, and D.~H. {Hartmann}.
\newblock {Formation Rates of Black Hole Accretion Disk Gamma-Ray Bursts}.
\newblock {\em \apj}, 526:152--177, Nov. 1999.

\bibitem{KN_magnetar}
H.~{Gao}, B.~{Zhang}, H.-J. {L{\"u}}, and Y.~{Li}.
\newblock {Searching for Magnetar-powered Merger-novae from Short GRBS}.
\newblock {\em \apj}, 837:50, Mar. 2017.

\bibitem{TeV_GRB_3}
H.~{Gao}, B.~{Zhang}, X.-F. {Wu}, and Z.-G. {Dai}.
\newblock {Possible high-energy neutrino and photon signals from gravitational
  wave bursts due to double neutron star mergers}.
\newblock {\em \prd}, 88(4):043010, Aug. 2013.

\bibitem{ghirlanda2019compact}
G.~Ghirlanda, O.~Salafia, Z.~Paragi, M.~Giroletti, J.~Yang, B.~Marcote,
  J.~Blanchard, I.~Agudo, T.~An, M.~Bernardini, et~al.
\newblock Compact radio emission indicates a structured jet was produced by a
  binary neutron star merger.
\newblock {\em Science}, page eaau8815, 2019.

\bibitem{GW170817_GBM}
A.~Goldstein, P.~Veres, E.~Burns, M.~Briggs, R.~Hamburg, D.~Kocevski,
  C.~Wilson-Hodge, R.~Preece, S.~Poolakkil, O.~Roberts, et~al.
\newblock An ordinary short gamma-ray burst with extraordinary implications:
  Fermi-gbm detection of grb 170817a.
\newblock {\em The Astrophysical Journal Letters}, 848(2):L14, 2017.

\bibitem{kilonova_diversity_observed}
B.~Gompertz, A.~Levan, N.~Tanvir, J.~Hjorth, S.~Covino, P.~Evans, A.~Fruchter,
  C.~Gonz{\'a}lez-Fern{\'a}ndez, Z.~Jin, J.~Lyman, et~al.
\newblock The diversity of kilonova emission in short gamma-ray bursts.
\newblock {\em The Astrophysical Journal}, 860(1):62, 2018.

\bibitem{guidorzi_2017_H0}
C.~Guidorzi, R.~Margutti, D.~Brout, D.~Scolnic, W.~Fong, K.~Alexander,
  P.~Cowperthwaite, J.~Annis, E.~Berger, P.~Blanchard, et~al.
\newblock Improved constraints on h 0 from a combined analysis of
  gravitational-wave and electromagnetic emission from gw170817.
\newblock {\em The Astrophysical Journal Letters}, 851(2):L36, 2017.

\bibitem{GW170817_afterglow_discovery_radio}
G.~Hallinan, A.~Corsi, K.~Mooley, K.~Hotokezaka, E.~Nakar, M.~Kasliwal,
  D.~Kaplan, D.~Frail, S.~Myers, T.~Murphy, et~al.
\newblock A radio counterpart to a neutron star merger.
\newblock {\em Science}, page eaap9855, 2017.

\bibitem{FRIB_rprocess}
C.~Horowitz, A.~Arcones, B.~C{\^o}t{\'e}, I.~Dillmann, W.~Nazarewicz,
  I.~Roederer, H.~Schatz, A.~Aprahamian, D.~Atanasov, A.~Bauswein, et~al.
\newblock r-process nucleosynthesis: Connecting rare-isotope beam facilities
  with the cosmos.
\newblock {\em arXiv preprint arXiv:1805.04637}, 2018.

\bibitem{hotokezaka_2018_H0}
K.~Hotokezaka, E.~Nakar, O.~Gottlieb, S.~Nissanke, K.~Masuda, G.~Hallinan,
  K.~P. Mooley, A.~Deller, et~al.
\newblock A hubble constant measurement from superluminal motion of the jet in
  gw170817.
\newblock {\em arXiv preprint arXiv:1806.10596}, 2018.

\bibitem{IceCube-Gen2}
{IceCube-Gen2 Collaboration}, {:}, M.~G. {Aartsen}, M.~{Ackermann}, J.~{Adams},
  J.~A. {Aguilar}, M.~{Ahlers}, M.~{Ahrens}, D.~{Altmann}, T.~{Anderson}, and
  et~al.
\newblock {IceCube-Gen2: A Vision for the Future of Neutrino Astronomy in
  Antarctica}.
\newblock {\em arXiv e-prints}, Dec. 2014.

\bibitem{KN_050709}
Z.-P. {Jin}, K.~{Hotokezaka}, X.~{Li}, M.~{Tanaka}, P.~{D'Avanzo}, Y.-Z. {Fan},
  S.~{Covino}, D.-M. {Wei}, and T.~{Piran}.
\newblock {The Macronova in GRB 050709 and the GRB-macronova connection}.
\newblock {\em Nature Communications}, 7:12898, Sept. 2016.

\bibitem{GW170817_kilonova_theory}
D.~Kasen, B.~Metzger, J.~Barnes, E.~Quataert, and E.~Ramirez-Ruiz.
\newblock Origin of the heavy elements in binary neutron-star mergers from a
  gravitational-wave event.
\newblock {\em Nature}, 551(7678):80, 2017.

\bibitem{GW170817_cocoon_Kasliwal}
M.~Kasliwal, E.~Nakar, L.~Singer, D.~Kaplan, D.~Cook, A.~Van~Sistine, R.~Lau,
  C.~Fremling, O.~Gottlieb, J.~Jencson, et~al.
\newblock Illuminating gravitational waves: a concordant picture of photons
  from a neutron star merger.
\newblock {\em Science}, 358(6370):1559--1565, 2017.

\bibitem{TeV_GRB_2}
S.~S. {Kimura}, K.~{Murase}, I.~{Bartos}, K.~{Ioka}, I.~S. {Heng}, and
  P.~{M{\'e}sz{\'a}ros}.
\newblock {Transejecta high-energy neutrino emission from binary neutron star
  mergers}.
\newblock {\em \prd}, 98(4):043020, Aug. 2018.

\bibitem{UHECR_A}
S.~S. Kimura, K.~Murase, and P.~M{\'e}sz{\'a}ros.
\newblock Super-knee cosmic rays from galactic neutron star merger remnants.
\newblock {\em The Astrophysical Journal}, 866(1):51, 2018.

\bibitem{TeV_GRB_1}
S.~S. {Kimura}, K.~{Murase}, P.~{M{\'e}sz{\'a}ros}, and K.~{Kiuchi}.
\newblock {High-energy Neutrino Emission from Short Gamma-Ray Bursts: Prospects
  for Coincident Detection with Gravitational Waves}.
\newblock {\em \apjl}, 848:L4, Oct. 2017.

\bibitem{GRBs_synchrotron}
P.~Kumar and B.~Zhang.
\newblock The physics of gamma-ray bursts and relativistic jets.
\newblock {\em arXiv preprint arXiv:1410.0679}, 2014.

\bibitem{MeV_neutrino_2}
K.~Kyutoku and K.~Kashiyama.
\newblock Detectability of thermal neutrinos from binary neutron-star mergers
  and implications for neutrino physics.
\newblock {\em Physical Review D}, 97(10):103001, 2018.

\bibitem{NS_EOS_review_Lattimer}
J.~Lattimer and M.~Prakash.
\newblock Neutron star structure and the equation of state.
\newblock {\em The Astrophysical Journal}, 550(1):426, 2001.

\bibitem{NSBH_rprocess_Lattimer}
J.~M. {Lattimer} and D.~N. {Schramm}.
\newblock {Black-hole-neutron-star collisions}.
\newblock {\em \apjl}, 192:L145--L147, Sept. 1974.

\bibitem{jet_form_2}
W.-H. {Lei}, B.~{Zhang}, X.-F. {Wu}, and E.-W. {Liang}.
\newblock {Hyperaccreting Black Hole as Gamma-Ray Burst Central Engine. II.
  Temporal Evolution of the Central Engine Parameters during the Prompt and
  Afterglow Phases}.
\newblock {\em \apj}, 849:47, Nov. 2017.

\bibitem{EH_Eichler}
A.~Levinson and D.~Eichler.
\newblock Baryon purity in cosmological gamma-ray bursts as a manifestation of
  event horizons.
\newblock {\em The Astrophysical Journal}, 418:386, 1993.

\bibitem{Li_1998}
L.-X. {Li} and B.~{Paczy{\'n}ski}.
\newblock {Transient Events from Neutron Star Mergers}.
\newblock {\em \apjl}, 507:L59--L62, Nov. 1998.

\bibitem{SGRB_magnetar_lu_2015}
H.-J. {L{\"u}}, B.~{Zhang}, W.-H. {Lei}, Y.~{Li}, and P.~D. {Lasky}.
\newblock {The Millisecond Magnetar Central Engine in Short GRBs}.
\newblock {\em \apj}, 805:89, June 2015.

\bibitem{GW170817_afterglow_optical_discovery}
J.~Lyman, G.~Lamb, A.~Levan, I.~Mandel, N.~Tanvir, S.~Kobayashi, B.~Gompertz,
  J.~Hjorth, A.~Fruchter, T.~Kangas, et~al.
\newblock The optical afterglow of the short gamma-ray burst associated with
  gw170817.
\newblock {\em arXiv preprint arXiv:1801.02669}, 2018.

\bibitem{NS_EOS_margalit_2017}
B.~Margalit and B.~D. Metzger.
\newblock Constraining the maximum mass of neutron stars from multi-messenger
  observations of gw170817.
\newblock {\em The Astrophysical Journal Letters}, 850(2):L19, 2017.

\bibitem{GW170817_structured_jet_margutti}
R.~Margutti, E.~Berger, W.~Fong, C.~Guidorzi, K.~Alexander, B.~Metzger,
  P.~Blanchard, P.~Cowperthwaite, R.~Chornock, T.~Eftekhari, et~al.
\newblock The electromagnetic counterpart of the binary neutron star merger
  ligo/virgo gw170817. v. rising x-ray emission from an off-axis jet.
\newblock {\em The Astrophysical Journal Letters}, 848(2):L20, 2017.

\bibitem{metzger2010electromagnetic}
B.~Metzger, G.~Mart{\'\i}nez-Pinedo, S.~Darbha, E.~Quataert, A.~Arcones,
  D.~Kasen, R.~Thomas, P.~Nugent, I.~Panov, and N.~T. Zinner.
\newblock Electromagnetic counterparts of compact object mergers powered by the
  radioactive decay of r-process nuclei.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  406(4):2650--2662, 2010.

\bibitem{SGRB_magnetar_metzger_2008}
B.~Metzger, E.~Quataert, and T.~A. Thompson.
\newblock Short-duration gamma-ray bursts with extended emission from
  protomagnetar spin-down.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  385(3):1455--1460, 2008.

\bibitem{review_grb_central_engine}
B.~D. Metzger.
\newblock Gamma-ray burst central engines: Black hole vs. magnetar.
\newblock {\em arXiv preprint arXiv:1001.5046}, 2010.

\bibitem{review_kilonovae_metzger}
B.~D. Metzger.
\newblock Kilonovae.
\newblock {\em Living Reviews in Relativity}, 20(1):3, 2017.

\bibitem{NS_EOS_metzger}
B.~D. Metzger and R.~Fern{\'a}ndez.
\newblock Red or blue? a potential kilonova imprint of the delay until black
  hole formation following a neutron star merger.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  441(4):3444--3453, 2014.

\bibitem{metzger2014optical}
B.~D. Metzger and A.~L. Piro.
\newblock Optical and x-ray emission from stable millisecond magnetars formed
  from the merger of binary neutron stars.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  439(4):3916--3930, 2014.

\bibitem{GW170817_VLBI_Mooley}
K.~Mooley, A.~Deller, O.~Gottlieb, E.~Nakar, G.~Hallinan, S.~Bourke, D.~Frail,
  A.~Horesh, A.~Corsi, and K.~Hotokezaka.
\newblock Superluminal motion of a relativistic jet in the neutron-star merger
  gw170817.
\newblock {\em Nature}, 561(7723):355, 2018.

\bibitem{GW170817_cocoon_Mooley}
K.~Mooley, E.~Nakar, K.~Hotokezaka, G.~Hallinan, A.~Corsi, D.~Frail, A.~Horesh,
  T.~Murphy, E.~Lenc, D.~Kaplan, et~al.
\newblock A mildly relativistic wide-angle outflow in the neutron-star merger
  event gw170817.
\newblock {\em Nature}, 554(7691):207, 2018.

\bibitem{NS_EOS_most_2018}
E.~R. Most, L.~J. Papenfort, V.~Dexheimer, M.~Hanauske, S.~Schramm,
  H.~St{\"o}cker, and L.~Rezzolla.
\newblock Signatures of quark-hadron phase transitions in general-relativistic
  neutron-star mergers.
\newblock {\em arXiv preprint arXiv:1807.03684}, 2018.

\bibitem{KM_2}
K.~Murase.
\newblock Prompt high-energy neutrinos from gamma-ray bursts in photospheric
  and synchrotron self-compton scenarios.
\newblock {\em Physical Review D}, 78(10):101302, 2008.

\bibitem{nonthermal_Murase}
K.~{Murase}, M.~W. {Toomey}, K.~{Fang}, F.~{Oikonomou}, S.~S. {Kimura},
  K.~{Hotokezaka}, K.~{Kashiyama}, K.~{Ioka}, and P.~{M{\'e}sz{\'a}ros}.
\newblock {Double Neutron Star Mergers and Short Gamma-ray Bursts: Long-lasting
  High-energy Signatures and Remnant Dichotomy}.
\newblock {\em \apj}, 854:60, Feb. 2018.

\bibitem{TeV_add_3}
K.~Murase, B.~Zhang, K.~Takahashi, and S.~Nagataki.
\newblock Possible effects of pair echoes on gamma-ray burst afterglow
  emission.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  396(4):1825--1832, 2009.

\bibitem{Narayan1992}
R.~{Narayan}, B.~{Paczynski}, and T.~{Piran}.
\newblock {Gamma-ray bursts as the death throes of massive binary stars}.
\newblock {\em \apjl}, 395:L83--L86, Aug. 1992.

\bibitem{SGRB_EE}
J.~P. Norris and J.~T. Bonnell.
\newblock Short gamma-ray bursts with extended emission.
\newblock {\em The Astrophysical Journal}, 643(1):266, 2006.

\bibitem{xray_afterglow_fading}
M.~{Nynka}, J.~J. {Ruan}, D.~{Haggard}, and P.~A. {Evans}.
\newblock {Fading of the X-Ray Afterglow of Neutron Star Merger GW170817/GRB
  170817A at 260 Days}.
\newblock {\em \apjl}, 862:L19, Aug. 2018.

\bibitem{NS_EOS_review_Ozel}
F.~{\"O}zel and P.~Freire.
\newblock Masses, radii, and the equation of state of neutron stars.
\newblock {\em Annual Review of Astronomy and Astrophysics}, 54:401--440, 2016.

\bibitem{Paczynski1991}
B.~{Paczynski}.
\newblock {Cosmological gamma-ray bursts}.
\newblock {\em Acta Astronomica}, 41:257--267, 1991.

\bibitem{GRB_afterglow_synchrotron}
T.~Piran, R.~Narayan, et~al.
\newblock Spectra and light curves of gamma-ray burst afterglows.
\newblock {\em The Astrophysical Journal Letters}, 497(1):L17, 1998.

\bibitem{Popham_1999}
R.~{Popham}, S.~E. {Woosley}, and C.~{Fryer}.
\newblock {Hyperaccreting Black Holes and Gamma-Ray Bursts}.
\newblock {\em \apj}, 518:356--374, June 1999.

\bibitem{qian_2000_summary}
Y.-Z. Qian.
\newblock Supernovae versus neutron star mergers as the major r-process
  sources.
\newblock {\em The Astrophysical Journal Letters}, 534(1):L67, 2000.

\bibitem{NS_EOS_radice_2018}
D.~Radice and L.~Dai.
\newblock Multimessenger parameter estimation of gw170817.
\newblock {\em arXiv preprint arXiv:1810.12917}, 2018.

\bibitem{NS_EOS_Radice_2017}
D.~Radice, A.~Perego, F.~Zappa, and S.~Bernuzzi.
\newblock Gw170817: Joint constraint on the neutron star equation of state from
  multimessenger observations.
\newblock {\em The Astrophysical Journal Letters}, 852(2):L29, 2018.

\bibitem{H0-Riess_2018}
A.~G. Riess, S.~Casertano, W.~Yuan, L.~Macri, B.~Bucciarelli, M.~G. Lattanzi,
  J.~W. MacKenty, J.~B. Bowers, W.~Zheng, A.~V. Filippenko, et~al.
\newblock Milky way cepheid standards for measuring cosmic distances and
  application to gaia dr2: Implications for the hubble constant.
\newblock {\em The Astrophysical Journal}, 861(2):126, 2018.

\bibitem{Riess_H0_38}
A.~G. Riess, S.~Casertano, W.~Yuan, L.~Macri, B.~Bucciarelli, M.~G. Lattanzi,
  J.~W. MacKenty, J.~B. Bowers, W.~Zheng, A.~V. Filippenko, et~al.
\newblock Milky way cepheid standards for measuring cosmic distances and
  application to gaia dr2: implications for the hubble constant.
\newblock {\em The Astrophysical Journal}, 861(2):126, 2018.

\bibitem{H0_Riess_2016}
A.~G. Riess, L.~M. Macri, S.~L. Hoffmann, D.~Scolnic, S.~Casertano, A.~V.
  Filippenko, B.~E. Tucker, M.~J. Reid, D.~O. Jones, J.~M. Silverman, et~al.
\newblock A 2.4\% determination of the local value of the hubble constant.
\newblock {\em The Astrophysical Journal}, 826(1):56, 2016.

\bibitem{UHECR_B}
X.~Rodrigues, D.~Biehl, D.~Boncioli, and A.~M. Taylor.
\newblock Binary neutron star merger remnants as sources of cosmic rays below
  the “ankle”.
\newblock {\em Astroparticle Physics}, 106:10--17, 2019.

\bibitem{NSM_Rosswog_1998}
S.~{Rosswog}, M.~{Liebend{\"o}rfer}, F.-K. {Thielemann}, M.~B. {Davies},
  W.~{Benz}, and T.~{Piran}.
\newblock {Mass ejection in neutron star mergers}.
\newblock {\em \aap}, 341:499--526, Jan. 1999.

\bibitem{SGRB_magnetar_rowlinson_2013}
A.~Rowlinson, P.~O'brien, B.~Metzger, N.~Tanvir, and A.~Levan.
\newblock Signatures of magnetar central engines in short grb light curves.
\newblock {\em Monthly Notices of the Royal Astronomical Society},
  430(2):1061--1087, 2013.

\bibitem{jets_neutrino_antineutrino}
M.~{Ruffert} and H.~T. {Janka}.
\newblock {Colliding neutron stars. Gravitational waves, neutrino emission, and
  gamma-ray bursts}.
\newblock {\em \aap}, 338:535--555, Oct. 1998.

\bibitem{merger_rates_review_Sadowski}
A.~Sadowski, K.~Belczynski, T.~Bulik, N.~Ivanova, F.~A. Rasio, and
  R.~O’Shaughnessy.
\newblock The total merger rate of compact object binaries in the local
  universe.
\newblock {\em The Astrophysical Journal}, 676(2):1162, 2008.

\bibitem{GW_IFO_3rd_Gen_EU}
B.~Sathyaprakash, M.~Abernathy, F.~Acernese, P.~Ajith, B.~Allen,
  P.~Amaro-Seoane, N.~Andersson, S.~Aoudia, K.~Arun, P.~Astone, et~al.
\newblock Scientific objectives of einstein telescope.
\newblock {\em Classical and Quantum Gravity}, 29(12):124013, 2012.

\bibitem{GW170817_INTEGRAL}
V.~{Savchenko}, C.~{Ferrigno}, E.~{Kuulkers}, A.~{Bazzano}, E.~{Bozzo},
  S.~{Brandt}, J.~{Chenevez}, T.~J.-L. {Courvoisier}, R.~{Diehl}, A.~{Domingo},
  L.~{Hanlon}, E.~{Jourdain}, A.~{von Kienlin}, P.~{Laurent}, F.~{Lebrun},
  A.~{Lutovinov}, A.~{Martin-Carrillo}, S.~{Mereghetti}, L.~{Natalucci},
  J.~{Rodi}, J.-P. {Roques}, R.~{Sunyaev}, and P.~{Ubertini}.
\newblock {INTEGRAL Detection of the First Prompt Gamma-Ray Signal Coincident
  with the Gravitational-wave Event GW170817}.
\newblock {\em \apjl}, 848:L15, Oct. 2017.

\bibitem{H0-Schutz-1986}
B.~F. Schutz.
\newblock Determining the hubble constant from gravitational wave observations.
\newblock {\em Nature}, 323(6086):310, 1986.

\bibitem{schutz_2002}
B.~F. Schutz.
\newblock Lighthouses of gravitational wave astronomy.
\newblock In {\em Lighthouses of the Universe: The Most Luminous Celestial
  Objects and Their Use for Cosmology}, pages 207--224. Springer, 2002.

\bibitem{GW_inclination}
B.~F. Schutz.
\newblock Networks of gravitational wave detectors and three figures of merit.
\newblock {\em Classical and Quantum Gravity}, 28(12):125023, 2011.

\bibitem{MeV_neutrino_1}
Y.~Sekiguchi, K.~Kiuchi, K.~Kyutoku, and M.~Shibata.
\newblock Gravitational waves and neutrino emission from the merger of binary
  neutron stars.
\newblock {\em Physical review letters}, 107(5):051102, 2011.

\bibitem{KM_1}
I.~M. Shoemaker and K.~Murase.
\newblock Constraints from the time lag between gravitational waves and gamma
  rays: Implications of gw170817 and grb 170817a.
\newblock {\em Physical Review D}, 97(8):083013, 2018.

\bibitem{rprocess_1_siegel}
D.~M. Siegel, J.~Barnes, and B.~D. Metzger.
\newblock The neutron star merger gw170817 points to collapsars as the main
  r-process source.
\newblock {\em arXiv preprint arXiv:1810.00098}, 2018.

\bibitem{siegel2016electromagnetic}
D.~M. Siegel and R.~Ciolfi.
\newblock Electromagnetic emission from long-lived binary neutron star merger
  remnants. i. formulation of the problem.
\newblock {\em The Astrophysical Journal}, 819(1):14, 2016.

\bibitem{TeV_add_2}
H.~Takami, K.~Kyutoku, and K.~Ioka.
\newblock High-energy radiation from remnants of neutron star binary mergers.
\newblock {\em Physical Review D}, 89(6):063006, 2014.

\bibitem{KN_130603B}
N.~R. {Tanvir}, A.~J. {Levan}, A.~S. {Fruchter}, J.~{Hjorth}, R.~A. {Hounsell},
  K.~{Wiersema}, and R.~L. {Tunnicliffe}.
\newblock {A `kilonova' associated with the short-duration {$\gamma$}-ray burst
  GRB 130603B}.
\newblock {\em \nat}, 500:547--549, Aug. 2013.

\bibitem{GW170817_afterglow_Xray_discovery}
E.~Troja, L.~Piro, H.~van Eerten, R.~Wollaeger, M.~Im, O.~Fox, N.~Butler,
  S.~Cenko, T.~Sakamoto, C.~Fryer, et~al.
\newblock The x-ray counterpart to the gravitational-wave event.
\newblock {\em Nature}, 551(7678):71, 2017.

\bibitem{SGRB_precursors}
E.~Troja, S.~Rosswog, and N.~Gehrels.
\newblock Precursors of short gamma-ray bursts.
\newblock {\em The Astrophysical Journal}, 723(2):1711, 2010.

\bibitem{KN_150101B}
E.~{Troja}, G.~{Ryan}, L.~{Piro}, H.~{van Eerten}, S.~B. {Cenko}, Y.~{Yoon},
  S.-K. {Lee}, M.~{Im}, T.~{Sakamoto}, P.~{Gatkine}, A.~{Kutyrev}, and
  S.~{Veilleux}.
\newblock {A luminous blue kilonova and an off-axis jet from a compact binary
  merger at z = 0.1341}.
\newblock {\em Nature Communications}, 9:4089, Oct. 2018.

\bibitem{tsang2012resonant}
D.~Tsang, J.~S. Read, T.~Hinderer, A.~L. Piro, and R.~Bondarescu.
\newblock Resonant shattering of neutron star crusts.
\newblock {\em Physical Review Letters}, 108(1):011102, 2012.

\bibitem{GW170817_kilonova_infrared}
V.~Villar, P.~Cowperthwaite, E.~Berger, P.~Blanchard, S.~Gomez, K.~Alexander,
  R.~Margutti, R.~Chornock, T.~Eftekhari, G.~Fazio, et~al.
\newblock Spitzer space telescope infrared observations of the binary neutron
  star merger gw170817.
\newblock {\em arXiv preprint arXiv:1805.08192}, 2018.

\bibitem{jet_process_gw170817}
Y.-Z. Wang, D.-S. Shao, J.-L. Jiang, S.-P. Tang, X.-X. Ren, Z.-P. Jin, Y.-Z.
  Fan, and D.-M. Wei.
\newblock Gw170817: The energy extraction process of the off-axis relativistic
  outflow and the constraint on the equation of state of neutron stars.
\newblock {\em arXiv preprint arXiv:1811.02558}, 2018.

\bibitem{KM_3}
E.~Waxman and J.~Bahcall.
\newblock High energy neutrinos from cosmological gamma-ray burst fireballs.
\newblock {\em Physical Review Letters}, 78(12):2292, 1997.

\bibitem{WEP_2}
J.-J. {Wei}, B.-B. {Zhang}, X.-F. {Wu}, H.~{Gao}, P.~{M{\'e}sz{\'a}ros},
  B.~{Zhang}, Z.-G. {Dai}, S.-N. {Zhang}, and Z.-H. {Zhu}.
\newblock {Multimessenger tests of the weak equivalence principle from GW170817
  and its electromagnetic counterparts}.
\newblock {\em JCAP}, 11:035, Nov. 2017.

\bibitem{Review_Tests_of_GR_overall_Will_2014}
C.~M. Will.
\newblock The confrontation between general relativity and experiment.
\newblock {\em Living Reviews in Relativity}, 17(1):4, 2014.

\bibitem{joint_williamson}
A.~R. Williamson, C.~Biwer, S.~Fairhurst, I.~Harry, E.~Macdonald, D.~Macleod,
  and V.~Predoi.
\newblock Improved methods for detecting gravitational waves associated with
  short gamma-ray bursts.
\newblock {\em Physical Review D}, 90(12):122004, 2014.

\bibitem{KN_060614}
B.~{Yang}, Z.-P. {Jin}, X.~{Li}, S.~{Covino}, X.-Z. {Zheng}, K.~{Hotokezaka},
  Y.-Z. {Fan}, T.~{Piran}, and D.-M. {Wei}.
\newblock {A possible macronova in the late afterglow of the long-short burst
  GRB 060614}.
\newblock {\em Nature Communications}, 6:7323, June 2015.

\bibitem{Zhang_Kumar}
B.~{Zhang} and P.~{Kumar}.
\newblock {Model-Dependent High-Energy Neutrino Flux from Gamma-Ray Bursts}.
\newblock {\em Physical Review Letters}, 110(12):121101, Mar. 2013.

\bibitem{GRB170817A_prompt_example}
B.-B. Zhang, B.~Zhang, H.~Sun, W.-H. Lei, H.~Gao, Y.~Li, L.~Shao, Y.~Zhao,
  Y.-D. Hu, H.-J. L{\"u}, et~al.
\newblock A peculiar low-luminosity short gamma-ray burst from a double neutron
  star merger progenitor.
\newblock {\em Nature communications}, 9(1):447, 2018.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{times}
\usepackage{geometry}
\usepackage[utf8]{inputenc}
\usepackage{enumitem,amssymb}
\usepackage{ragged2e}
\usepackage{fancyhdr}
\usepackage{graphics,graphicx}
\usepackage{color}
\usepackage{xcolor}
\usepackage{lineno}
\usepackage{geometry}
\usepackage{wasysym}

\usepackage[sort&compress,numbers]{natbib}

\usepackage[compact]{titlesec}

\newlist{thematic}{itemize}{8}
\setlist[thematic]{label=$\square$}
\usepackage{pifont}
\newcommand{\cmark}{\ding{51}}%
\newcommand{\done}{\rlap{$\square$}{\raisebox{2pt}{\large\hspace{1pt}\cmark}}%
\hspace{-2.5pt}}

\begin{document}
\raggedright
\huge
Astro2020 Science White Paper \linebreak

A Summary of Multimessenger Science with Neutron Star Mergers \linebreak
\normalsize

\noindent \textbf{Thematic Areas:} \hspace*{60pt} $\square$ Planetary Systems \hspace*{10pt} $\square$ Star and Planet Formation \hspace*{20pt}\linebreak
$\done$ Formation and Evolution of Compact Objects \hspace*{31pt} $\done$ Cosmology and Fundamental Physics \linebreak
  $\square$  Stars and Stellar Evolution \hspace*{1pt} $\square$ Resolved Stellar Populations and their Environments \hspace*{40pt} \linebreak
  $\square$    Galaxy Evolution   \hspace*{45pt} $\done$             Multi-Messenger Astronomy and Astrophysics \hspace*{65pt} \linebreak
  
\textbf{Principal Author:}

Name:	Eric Burns
 \linebreak						
Institution:  NASA Goddard
 \linebreak
Email: eric.burns@nasa.gov
 \linebreak
Phone:  +1-301-286-4664
 \linebreak

\textbf{Co-authors:} 

A. Tohuvavohu (Pennsylvania State University), J. Buckley (Washington University), T. Dal Canton (NASA Goddard), S. B. Cenko (NASA Goddard), J. W. Conklin (University of Florida), F. D'Ammando (INAF - Istituto di Radioastronomia), D. Eichler (Ben-Gurion University), C. Fryer (Los Alamos National Laboratory), A. J. van der Horst (George Washington University), M. Kamionkowski (Johns Hopkins), M. Kasliwal (Caltech), R. Margutti (Northwestern University), B. D. Metzger (Columbia University), K. Murase (Pennsylvania State University), S. Nissanke (GRAPPA), D. Radice (Princeton), J. Tomsick (UC Berkeley),  C. A. Wilson-Hodge (NASA Marshall), B. Zhang (University of Nevada, Las Vegas)
  \linebreak
  
\justify
\textbf{Abstract:}

Neutron star mergers, referring to both binary neutron star and neutron star black hole mergers, are the canonical multimessenger events. They have been detected across the electromagnetic spectrum, have recently been detected in gravitational waves, and are likely to produce neutrinos over several decades in energy. The non-thermal prompt and afterglow emission of short gamma-ray bursts and the quasi-thermal emission from the radioactively powered kilonovae provide distinct insights into the physics of neutron star mergers. When combined with direct information on coalescence from gravitational waves and neutrinos these sources may become the best understood astrophysical transients. Multimessenger observations of these cataclysmic events will determine sources of gravitational waves and astrophysical neutrinos, enable precision cosmology, and unique tests of fundamental physics, the origin of heavy elements, the behavior of relativistic jets, and the equation of state of supranuclear matter. In this white paper we present a summary of the science discoveries possible with multimessenger observations of neutron star mergers and provide recommendations to enable them in the new era of time-domain, multimessenger astronomy.

\pagebreak

\section{Introduction}
The first multimessenger detection of a neutron star (NS) merger was jointly detected in gravitational waves (GWs) and $\gamma$-rays \citep{GW170817_LVC,GW170817_GBM,GW170817_INTEGRAL,GW170817-GRB170817A}. The aftermath was detected in follow-up observations across the electromagnetic (EM) spectrum \citep[e.g][]{GW170817_afterglow_Xray_discovery,GW170817_afterglow_discovery_radio,GW170817_afterglow_optical_discovery, xray_afterglow_fading,GW170817_kilonova_optical,GW170817_kilonova_UV,GW170817_kilonova_infrared}. We describe the outstanding science possible with observations of NS mergers, followed by the capabilities necessary for discovery. 

\section{Science}
This science is predicated on astrophysical observations of these phenomena; therefore, we first discuss the astrophysics of NS mergers and follow with the broader science they enable. 

\subsection{Astrophysics}
The detections and associations of GW170817, GRB 170817A and the kilonova, AT2017gfo, confirmed decades old predictions from theory, observation, and simulation on the nature of short gamma-ray bursts (SGRBs) and the origin of heavy (r-process) elements \citep[e.g.][]{NSBH_rprocess_Lattimer, Blinnikov1984, Eichler_1989, Paczynski1991, Narayan1992, jets_neutrino_antineutrino, NSM_Rosswog_1998, Li_1998, NSM_Bloom_1999, Popham_1999, Fryer_1999, SGRBs_fong}. We first discuss the (mostly) distinct SGRB and kilonovae transients, followed by astrophysical highlights informed by studies of these sources. 

\subsubsection{Short Gamma-Ray Bursts}
Gamma-Ray Bursts (GRBs) originate from collimated, relativistic outflows. Their prompt emission are the most luminous (EM) events in the universe, but we do not understand where or how the $\gamma$-rays are produced \citep[see, e.g.][for reviews]{GRBs_synchrotron,GRBs_photospheric}. Prompt emission is followed by afterglow from synchrotron radiation as the jet interacts with surrounding material \citep{GRB_afterglow_synchrotron}, observations of which enable studies of the local environment and energetic efficiencies \citep[e.g.][]{SGRBs_fong}. GRB 170817A was underluminous compared to the existing sample of SGRBs, and the temporal behavior of its afterglow had never been previously observed \citep[e.g.][]{GW170817_GBM,GW170817-GRB170817A,GW170817_afterglow_Xray_discovery,GW170817_structured_jet_margutti}. This led to significant debate on the structure of the non-thermal emitting region \citep[e.g.][]{GW170817_kilonova_UV,GW170817_afterglow_Xray_discovery,GW170817_structured_jet_margutti, GW170817_cocoon_Mooley,GRB170817A_Eichler,GW170817_afterglow_latettime_2018,GW170817_VLBI_Mooley,ghirlanda2019compact} and the origin of the prompt emission \citep[e.g.][]{GW170817_cocoon_Kasliwal,GW170817-GRB170817A,GRB170817A_prompt_example,KM_1}.

Multimessenger studies of NS mergers will determine the origin and implications of their non-thermal emission (prompt, afterglow, and more) \citep[e.g.][]{GW170817-GRB170817A,GRB170817A_prompt_example,SGRB_precursors,SGRB_EE,SGRB_magnetar_rowlinson_2013,SGRB_precursors,tsang2012resonant,nonthermal_Murase} and their relation to the different possible central engines \citep[e.g.][]{SGRB_magnetar_lu_2015, SGRB_magnetar_metzger_2008}, the structure of SGRB jets and their interaction with the kilonova material \citep[e.g.][]{GW170817_kilonova_UV}, and how they relate to the intrinsic parameters of the progenitor system. These studies may soon be augmented by TeV detections of SGRBs \citep{TeV_GRB_1,TeV_GRB_2,CTA_1,CTA_2,TeV_add_2,TeV_add_3}. Well characterized events will uncover the fraction of SGRBs from binary neutron star (BNS) and neutron star black hole (NSBH) mergers, and the differences between the two. The merger time from GWs and the onset time of $\gamma$-rays in joint GW-GRB detections will constrain the bulk Lorentz factor of the jet (which determines the observable region of the jet due to Doppler beaming) and the size of the $\gamma$-ray emitting region \citep[as in][]{GW170817-GRB170817A}, and provide new information to uncover the origin of the prompt emission \citep[e.g.][]{GW170817-GRB170817A,GRB170817A_prompt_example}. 

\subsubsection{Kilonovae}
Prior searches for kilonovae following well-localized SGRBs resulted in a few claims of detection \citep{KN_130603B,KN_060614,KN_050709,KN_magnetar,KN_150101B} and interesting non-detections \citep[e.g.][]{kilonova_diversity_observed}; the only unambiguous, well-studied kilonova is the one that followed GW170817: AT2017gfo. This event met broad expectations, and a general picture has emerged: a bright blue kilonova from the relatively lanthanide-poor material in the polar regions followed by a transition to a fainter, progressively redder kilonova from lanthanide-rich material in the planar region \citep[e.g.][]{GW170817_kilonova_theory}. However, the  ultraviolet emission at 12 hours post merger was brighter than some expectations \citep[e.g.][]{metzger2010electromagnetic,GW170817_kilonova_UV} and earlier optical and ultraviolet observations are key to determining the origin of this emission \citep{GW170817_kilonova_models_early}. 

Detections of kilonovae following GW triggers will reveal the relationship between kilonova color and luminosity to properties of the progenitor, including merger type (BNS or NSBH), the short-term remnant object (BH, metastable NS, stable NS), the mass and spin of the progenitors, constrain the NS equation of state (EOS), and ultimately tie the observed diversity to specific physical processes \citep[see][for a review]{review_kilonovae_metzger}.

\subsubsection{Neutron Star-Black Hole Systems}
NSBH mergers are also predicted to result in both SGRBs and kilonova, so long as the NS is disrupted outside of the innermost stable circular orbit. However, there is no known stellar system that is definitely a NSBH binary. Since NSBH mergers have larger detectable volumes than BNS mergers \citep[e.g.][]{LVCK_prospects} we should expect a detection in the next few years, unless they are significantly less common. The first classification of a NSBH merger will likely require multimessenger observations (unless the GW signal is particularly loud). The BH can be inferred from the GW mass measures and the NS from an associated SGRB or predominantly red kilonova. Such an observation would prove NSBH systems exist by observing one merge.

\subsubsection{Stellar Formation and Evolution}
Population synthesis studies bring together models of stellar formation, stellar evolution, binary evolution, and supernovae to make predictions on the properties of compact binaries. The formation channels of compact binary systems that merge in a Hubble time include the evolution of field binaries \citep[see][for a review]{BNS_review_rasio, dominik_2013} or dynamical capture. Multimessenger observations can classify GW triggers as NS mergers, measure mass and spin distributions, the offset from their host galaxies, the local volumetric merger rates of these systems and the source evolution of those rates. These observations with population synthesis studies can differentiate between formation channels, constrain the fraction of stellar systems that result in NS mergers \citep[see, e.g.][for results from GW170817]{GW170817_LVC_progenitor}, the initial offset distribution for the compact objects, the fraction of supernovae that result in NS or BH remnants and this relation to progenitor mass \citep[see, e.g.][]{merger_rates_review_Sadowski}. 

\subsection{Cosmology}
$H_0$ sets the age and size of the universe. It is one of six parameters in the base $\Lambda$CDM concordance cosmological model. It can be measured directly in the nearby (late) universe from type Ia supernovae studies \citep[e.g.][]{H0-Riess_2018}, or inferred from observations of the distant (early) universe from studies using observations of the Cosmic Microwave Background (CMB), Baryon Acoustic Oscillations (BAO), and an assumed cosmological model \citep[e.g.][]{Planck_2018_cosmo}. Comparing $H_0$ values from observations of the early universe against those from observations of the late universe provides a stringent test of cosmological models. These measures disagree with 3.8$\sigma$ significance \citep{Riess_H0_38} with no obvious systematic origin \citep{H0_Riess_2016,Planck_2018_cosmo}. 

Mergers of compact objects are standard sirens with an intrinsic luminosity predicted from GR which can be combined with an associated redshift to measure $H_0$ \citep{H0-Schutz-1986,schutz_2002}. The precision is limited by the correlated inclination and distance determination from GW measures, which can be ameliorated for events with an associated SGRB \citep[e.g.][]{dalal_2006,guidorzi_2017_H0,GW170817_VLBI_Mooley,hotokezaka_2018_H0}. Over the next decade standard siren cosmology could resolve the current $H_0$ controversy and independently calibrate the cosmological ladder \citep[see][for a review]{freedman2001final}. Further, as the GW interferometers improve and detect NS mergers deep into the universe, those with (EM-determined) redshift will create the most precise Hubble Diagram spanning from the local universe to deep into the deceleration era \citep{GW_IFO_3rd_Gen_EU}. Combining standard siren cosmology with CMB+BAO observations could resolve the neutrino mass hierarchy question and improve constraints on the effective number of neutrino species and the EOS of dark energy \citep{Planck_2018_cosmo}. Further adding information from forthcoming transition era cosmology experiments (e.g. LSST, EUCLID, WFIRST) gives sub-percent precision cosmology throughout the universe, allowing studies on multi-parameter extensions to $\Lambda$CDM (e.g. $\Omega_k$, $w_0$, and $w_a$ simultaneously). 

\subsection{Fundamental Physics}
The $\sim$seconds difference in the arrival times of gravity and light across cosmological baselines enable observations of NS mergers to probe some fundamental aspects of physics far greater than any other method. Detections with GWs and GRBs in $\sim$keV-MeV energy (which do not undergo significant extinction, dispersion, nor absorption) provide the best measure of the speed of gravity \citep{GW170817-GRB170817A}, and relative violations of the Weak Equivalence Principle \citep[e.g.][]{Review_Tests_of_GR_overall_Will_2014,GW170817-GRB170817A, WEP_2,KM_1} and Lorentz Invariance (LIV) \citep{LIV_SME,GW170817-GRB170817A,KM_1}. GeV-TeV (EM) observations of SGRBs have the greatest discovery space, and current best constraints, on general LIV \citep{LIV_090510_limit_2009}. These observations test the Special Theory of Relativity which has been woven into all of modern physics, the Einstein Equivalence Principle that all metric theories of gravity must obey, the quantization of spacetime itself (and therefore quantum gravity), and measure fundamental constants of the universe \citep[see e.g.][]{2015CQGra..32x3001B}. 

\subsection{The Origin of Heavy Elements}
For many decades, astronomers have debated whether NS mergers or Core-Collapse Supernovae (CCSN) are the site of heavy (above the iron peak), rapid neutron capture (r-process) elements \citep[see][for a summary of the arguments]{qian_2000_summary}. Even with the well-sampled AT2017gfo we are unsure if all three abundance peaks were synthesized. With a greater understanding of the r-process production in NS mergers from future kilonovae observations and improved measures on the rates of such events from GW observations, we can determine the total and relative heavy elemental abundances produced from these events. This can be compared to the observed solar system abundances to determine the fraction of heavy elements that come from NS mergers, as opposed to CCSN \citep[see e.g.][]{rprocess_1_siegel, rprocess_2_cote, rprocess_3_drout}. CCSN are thought to track the cosmic stellar formation rate, as are BNS and NSBH mergers modulo their inspiral times. An improved understanding of the source evolution of NS mergers will determine the heavy element abundance through cosmic time.

\subsection{Relativistic Jets and Particle Acceleration}
Many astrophysical sources, such as blazars, microquasars, and protostars, arise from collimated outflows referred to as jets; GRB jets are an ultrarelativistic version. Multimessenger observations provide direct measures of the central engine, which when combined with EM observations of the SGRB will uncover key information about jets. We will learn if SGRB jet formation requires an event horizon \citep{EH_Eichler} (i.e. a BH central engine), or if they can be created by magnetar central engines \citep[see, e.g.][]{review_grb_central_engine,TeV_GRB_3}. Multimessenger observations can determine required energetics efficiencies that may delineate between models of jet formation \citep[e.g.][]{jets_blandford_znajek,jets_neutrino_antineutrino,jet_process_gw170817, jet_form_2, TeV_GRB_1}. The structure of GRB jets can be studied from observations of the non-thermal emission, and move our understanding beyond the on-axis top-hat jet models rejected by GRB 170817A. Relativistic jets may be either hadron or magnetically dominated, which could be determined from (non-)detections of neutrinos \citep[e.g.][]{Zhang_Kumar,TeV_GRB_1,TeV_GRB_2,KM_2,KM_3,UHECR_A} with sufficiently sensitive detectors. The origin of cosmic rays beyond the knee or second knee energies is unknown, and neutron star mergers have been suggested among the promising candidate sources \citep{UHECR_A, UHECR_B}.

\subsection{Neutron Star Equation of State}
NSs are the densest known matter in the universe, far beyond anything achievable in terrestrial laboratories. The NS EOS characterizes the density-pressure relationship \citep[see][for a review]{NS_EOS_review_Lattimer}. From an assumed EOS, several observables can be predicted, such as a mass-radius relation \citep[e.g.][]{NS_EOS_review_Ozel}. In BNS mergers the remnant object just after merger can be a stable NS, a metastable NS, or a BH, with the NS EOS determining which case for a given merger. These options can be resolved by observations of associated SGRBs or kilonovae with specific characteristics \citep[e.g.][]{SGRB_EE, SGRB_magnetar_metzger_2008, SGRB_magnetar_rowlinson_2013, metzger2014optical, siegel2016electromagnetic, nonthermal_Murase, GW170817-GRB170817A,NS_EOS_metzger,NS_EOS_Radice_2017, NS_EOS_margalit_2017, NS_EOS_radice_2018} and eventually directly confirmed with GW or MeV neutrino observations \citep{MeV_neutrino_1,MeV_neutrino_2}. Multimessenger observations of NSBH mergers provide additional constraints as the orbital radius for NS disruption is a strong function of the mass and spin of the BH and the NS radius \citep{NS_EOS_foucart_2012}. Multimessenger observations of NS mergers will provide information, complementary to EM-only or GW-only constraints, to determine the EOS of supranuclear matter, and possibly constrain the phase diagram of quantum chromodynamics \citep[e.g.][]{NS_EOS_most_2018, NS_EOS_bauswein_2018}. 

\section{Recommendations}
NS mergers emit across decades in energy in several messengers which enables greater understanding but requires vast observational resources. Prompt $\gamma$-ray, afterglow (TeV to radio), and TeV-PeV neutrino observations will provide new insight into SGRBs, relativistic jets, and particle acceleration. Ultraviolet, optical, and infrared (UVOIR) observations uncover and characterize kilonovae and the origin of heavy elements. Probing fundamental physics requires GW and prompt GRB detections. Cosmological studies require GW distance and EM redshift determinations. All observations help constrain the NS EOS. Much of this science is only possible with a population of events. Information on the inspiral, coalescence, or jet before the coasting phase (obtained from GW, $\gamma$-ray and neutrino detectors) can only be done from serendipitous observations, requiring all-sky monitors with high duty cycles and large fields of view.

Some detections also enable other observations. Despite collimation, joint GW-GRB detections are expected to be reasonably common because of preferential selection effects \citep[combining information from][]{GW_inclination,SGRBs_fong} and the increase of GW detections from GRB observations \citep[e.g.][]{joint_williamson, joint_blackburn}. Joint GW-GRB detections further constrain the region of interest for follow-up. Arcsecond localizations are necessary for host-galaxy identification and redshift determination and are only possible from follow-up observations. The earliest follow-up signal that can be detected is the GRB afterglow; however, based on the prompt vs afterglow brightness of GRB 170817A, we do not expect off-axis afterglow detections (with no detectable prompt emission) to be common. When there is no detectable SGRB, the first EM emission for a blue kilonova appears to be in the UV and in optical or infrared for red kilonovae (on longer timescales).

Kilonovae will be the dominant EM counterpart for nearby NS mergers as they are omnidirectional, but SGRBs will dominate for distant events as they can be significantly brighter. While we split our recommendations between the next decade and beyond, regardless of time period, \textit{we recommend vigorous funding for upgraded GW interferometers}.


\subsection{The Next Ten Years}
With the full design network of current GW detectors, localizations sufficient for existing optical facilities will be relatively common \citep{LVCK_prospects}. The funded LIGO A+ upgrade (nominally available in 2024), will detect a NS merger roughly once a week. Follow-up ground observations with current or expected missions (e.g. HAWC, CTA, VLA, SKA, ZTF, LSST) reliably cover TeV, optical, near-IR, and radio. As kilonovae transition to redder emission over time, the IR-only emission is sufficiently late ($\sim$days) that these wavelengths can be covered by JWST and WFIRST observations. \textit{We endorse the allocation of appropriate observing time and target of opportunity programs for pointed telescopes and directly recommend LSST follow-up of NS mergers}. Megaton-class MeV neutrino detectors could detect particularly nearby NS mergers. Initial upgrades to TeV-PeV neutrino detectors are on-going; \textit{we endorse the full IceCube Gen-2} \citep{IceCube-Gen2}. 

The critical wavelengths detectable only in space are keV-MeV $\gamma$-rays, X-rays, and UV. As such, we recommend the extension of the \textit{Fermi} mission (because it detects more prompt SGRBs than all other active mission combined) and the \textit{Neil Gehrels Swift Observatory} (primarily for fast-response X-ray and UV coverage). To capture the full range of possible kilonova colors, \textit{we recommend wide-field UV (space-based) and NIR (ground-based) facilities}. To enable time-domain astronomy, \textit{we recommend improvements to real-time communication for space-based missions.}

Beyond the required observational capabilities, \textit{we strongly recommend greater NSF-NASA cooperation} as both agencies have assets critical for multimessenger science. \textit{Specific recommendations include the funding of beam studies to understand the nuclear processes that currently limit kilonova models \citep[e.g.][]{review_kilonovae_metzger, FRIB_rprocess}, robust funding of multimessenger simulation and theory studies via the TCAN program (created in response to Astro2010), and the consideration of resolving grand problems through NASA-NSF partnership (akin to the DRIVE initiative from the 2013 Solar and Space Science Decadal).} \textit{Technical improvements include improvements to real-time reporting, automated multimission and multimessenger searches, and prompt reporting of initial parameter estimation from GW detections (e.g. masses) to enable follow-up prioritization.}

\subsection{Future Large-scale Missions}
The proposed LIGO-Voyager upgrade has a nominal timeline of $\sim$2030 and it would detect several NS mergers per week with some at cosmological distances ($z>0.1$). \textit{Proposed large-scale missions in the Astro2020 Decadal must be designed for the 2030s, not the current era.} Proposed third generation GW interferometers \citep{GW_IFO_3rd_Gen_US,GW_IFO_3rd_Gen_EU} with best-case timelines of mid-to-late 2030s could detect dozens of NS mergers per day with some beyond the transition era. Fewer upgraded interferometers will result in poorer GW localizations for distant events. To enable all NS merger science \textit{we recommend future GW interferometers aim for broadband sensitivity improvements}. With these upgrades a significant fraction of SGRBs will have associated GW emission. Therefore, \textit{we recommend a large-scale $\gamma$-ray observatory that will detect far more SGRBs than any prior mission (through improved $\sim$keV-MeV sensitivity and broad sky coverage) and localize them to sufficient accuracy for sensitive follow-up observations.} To enable a full study of these events, including redshift determination, they need to be localized to arcsecond precision. With such a $\gamma$-ray instrument, \textit{we would need well-matched follow-up facilities such as a sensitive, fast-reponse, high spatial resolution X-ray telescope and sensitive UVOIR and radio telescopes.}


\pagebreak

\bibliographystyle{abbrv}
\bibliography{references}

\end{document}